OpenStudy (anonymous):

find the exact value of the following? sin195

OpenStudy (anonymous):

\[\frac{ (-\sqrt{6} + \sqrt{2})}{ 4 }\]

OpenStudy (anonymous):

i got -1/2\[-1/2\sqrt{2-\sqrt{3}}\]

OpenStudy (anonymous):

\[\sqrt{(1-\cos(30))/2} = \sqrt{1/2(2/2 - \sqrt{3}/2}\]

OpenStudy (anonymous):

you use the half angle formula

OpenStudy (anonymous):

my teacher provided me with this sin20=cos^20(theta)*cos(theta) but i dont understand what do i do with it do i plug in the 195 instead?

OpenStudy (anonymous):

hmm you dont have any instructions? say something like prove the identity because sin2theta = 2sincos

OpenStudy (anonymous):

nope thats all that it says

OpenStudy (anonymous):

well the exact value of sin(195) is \[-1/2\sqrt{2-\sqrt{3}}\]

OpenStudy (anonymous):

wait nvm i read the half angle identity oops sorry

OpenStudy (anonymous):

using the half angle formula we double the angle value of 195 which comes to 390

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

since 390 is coterminal with 30 with use the value of cos(30) in the formula

OpenStudy (anonymous):

the formula is \[\sqrt{(1-\cos(\theta))/2}\]

OpenStudy (anonymous):

the cos of 30 is \[\sqrt{3/}2\]

OpenStudy (anonymous):

what do you mean by coterminal shouldnt the 30 be negative or something

OpenStudy (anonymous):

im just trying to understand it

OpenStudy (anonymous):

so the formula becomes \[\sqrt{(1-\sqrt{3}/2)/2}\]

OpenStudy (anonymous):

coterminal just means they are at the same spot on the unit circle

OpenStudy (anonymous):

since double 195 is 390 that means it goes all the way 360 degrees and then there is another 30 degrees

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

your angle went all the way around + 30 so its at pi/6

OpenStudy (anonymous):

alright i got that part

OpenStudy (anonymous):

so the cosine of 390 and 30 is the same

OpenStudy (anonymous):

since your angle is 195 we use the half angle formula

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

because 195 is hard to get an exact number for we double it which gives a good number that lands on 30degrees and just divide that by 2

OpenStudy (anonymous):

but the complete formula is \[\sqrt{(1-\cos(x))/2}\]

OpenStudy (anonymous):

wait so do i divide 30? by 2 or what do i divide?

OpenStudy (anonymous):

let me see if i can figure out how to put fractions into the equations im kinda new to openstudy

OpenStudy (anonymous):

ok thanks and i am to

OpenStudy (anonymous):

ill just draw it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):